OpenStudy (heart_offire):

What is the y-intercept of y = - 4 _ 3 x + 4? A) (0, 4) B) (4, 0) C) (0, -4) D) (-4, 0) @mathmale @neonumbrella5115 @study312

pooja195 (pooja195):

y=mx+b m=slope b=y intercept This is known as slope intercept form

pooja195 (pooja195):

Is your equation written like this: \[\huge~\rm~\bf~y=\frac{ 4 }{ 3 }x+4\]

OpenStudy (heart_offire):

its with a negative in front of the fraction

OpenStudy (neonumbrella5115):

If 4 is where the line hits the y intercept it would be located on the intercept, therefore having a 0 as its x value

pooja195 (pooja195):

\[\huge~\rm~\bf~y=-\frac{ 4 }{ 3 }x+4\]

OpenStudy (heart_offire):

but wouldnt the 0 just factor into the fraction

pooja195 (pooja195):

y=mx+b b=y intercept What is the y intercept in this equation?

OpenStudy (heart_offire):

@pooja195 yes thats what it looks like

pooja195 (pooja195):

Alright what is the y intercept?

OpenStudy (neonumbrella5115):

So the b is equivalent to the y value. What is the b? 4 or -4

pooja195 (pooja195):

I'll let @neonumbrella5115 help :-) don't want to confuse you.

OpenStudy (heart_offire):

does the x go in front of the ordered pair or in the back

OpenStudy (neonumbrella5115):

Coordinate pairs are in the form (x,y)

OpenStudy (heart_offire):

so it would be A

OpenStudy (neonumbrella5115):

Yes, you are correct!